# FaceQSORT: a Multi-Face Tracking Method based on Biometric and Appearance Features

### v2

@@ 2.1 Computational Complexity
As shown in Algorithm 1, FaceQSORT essentially comprises the matching cascade, the IoU fallback matching, the feature updating and the position prediction. In general, the complexity of finding bipartite matches using the Hungarian algorithm is $O(|\Phi|^{2}\times|\Psi|)$. This is also the worst case with cascade matching. In the considered scenario, it is very likely that the detected face $\Psi$ in frame $t$ were also present in frame $t-1$. Thus, it is reasonable that the complexity can be reduced by the matching cascade (i.e., to $O((\nicefrac{{|\Phi|}}{{x}})^{2}\times\Psi)$). For all unmatched detections after the matching cascade, the complexity of the IoU fallback matching is in worst case again $O(|\Phi|^{2}\times|\Psi|)$. The complexity for updating the memory for biometric and appearance features ($\Phi_{bio}$ and $\Phi_{app}$) is $O(|\Psi|)$ in each case. Finally, predicting the positions in frame $t+1$ using the Kalman filter has basically a complexity of $O(\Psi)$. Thus, the overall complexity of FaceQSORT is,
@@ 3 PLUS Faces in a Queue Dataset
The Paris Lodron University Salzburg Faces in a Queue (PLUSFiaQ) dataset is a new multi-face tracking dataset consisting of a total of seven different sequences. This dataset is made publicly available [^1] https://www.wavelab.at/sources/Joechl24b/. An overview of the individual sequences is given in Table 1. In total, the seven sequences comprise 12730 fully annotated frames. This corresponds to about 8 minutes and 30 seconds of video material (25 frames per second). The recorded sequences reflect the assumed scenario (described in section 1), in which several people move towards a gate in a queue. Informed consent was obtained from all participants for all released sequences.
@@ 3.1 Scenario Details
The assumed scenario (i.e., queue in front of a gate) is simulated with 12 different people. These persons move towards a camera equipped gate. On their way to the gate, they form a queue. However, this queue is not very well organized, i.e., people are chatting, eating, moving quickly, walking backwards, pushing or shoving. As soon as a person has entered the gate, they leave the visible area of the camera and thus the scene. To simulate multiple runs, each person goes more than once through the gate. For this reason, the people move in a counterclockwise circle (as depicted in Figure 2). Thus, when a person re-enters the scene, the person first moves away from the gate (camera). This is also illustrated in Figure 3, where the trajectory of the top-left Bounding-Box coordinates from a specific person is shown (i.e., trID 1702, marked by a red circle). At position (a) the person re-enters the scene, then moves away from the gate (b) until he is roughly farthest away at (c) and then moves towards the gate again at (d) and (e). Finally, the person again leaves the scene (goes through the gate) at position (f). This loop is repeated several times.
The only exception is sequence 1, in which only three people are involved, and they move clockwise, i.e., when they re-enter the scene, they move directly towards the gate. This sequence is also the most simple sequence.
@@ 3.2 Annotation Details
The trID is a four-digit number (e.g. xxyy), where the first two digits (xx) represent the personal identifiers and the last two digits (yy) represent a consecutive tracking number (per sequence). A new tracking number is assigned when a person has left the scene (goes through the gate) and re-enters the scene again. For example, the person in figure 3 (a), who is directly in front of the gate, has the personal identifier 18. The current tracking number is 1 (i.e., trID 1801). However, after leaving and re-entering the scene, the tracking number is incremented by one (i.e. trID 1802), as shown in Figure 3 (d). For all non-target class objects (faces) a personal identifier of 99 is assigned. The individual non-target class instances are assigned a dedicated tracking number, independent of how often they have re-entered the scene.
For annotation, the VGG Image Annotator (VIA) [^2] http://www.robots.ox.ac.uk/ vgg/software/via [50] was used. The corresponding exported annotation file (json file) is provided for each sequence. Furthermore, a Ground-Truth (GT) file according to the MOT20 format [51] is provided additionally, i.e., a csv file where each line represents one object instance and contains 9 values:
@@
As a starting point for the annotation, the BBs detected by yolov5333https://github.com/ultralytics/yolov5 and the trIDs predicted by StrongSORT [17] were used. These BBs and trIDs were then refined manually.
@@ 3.3 Tracking Challenges
4 Experimental Evaluation
FaceQSORT is evaluated based on the proposed PLUSFiaQ (see section 3) and the ChokePoint[47] dataset. In the ChokePoint dataset, people moving through a portal (chokepoint) are recorded. The dataset comprises sequences from 18 different runs (i.e., different portal, session or walking direction), with 3 different camera perspectives available for each run (i.e., $18\times 3=54$ available sequences). Of these 54 sequences, only 6 sequences show multiple persons per frame, which is similar to the assumed scenario in this paper. These 6 sequences are therefore used for evaluation. However, no annotations are available for these 6 sequences and are created accordingly. In the ChokePoint dataset, people are reflected in the open glass door as they walk through the portal. These refelections are annotated as non-target classes. The annotations created have the same format as described in section 3 and are made available along with the PLUSFiaQ dataset.
@@ 4.1 Tracker Implementation
FaceQSORT is implemented based on the StrongSORT implementation [^4] https://github.com/dyhBUPT/StrongSORT. The parameter $\alpha$ for the EMA feature updating strategy is set to $\alpha=0.9$. The weighting parameter $\beta$ (for weighting the contribution of biometric and appearance similarity compared to spatial distance) is set to 0.98. This gives the biometric and appearance similarity the most weight. $\theta_{pos}$ is set to 9.4877 (the 0.95 quantile of the chi-squared distribution with 4 degrees of freedom). These parameters are set according to [17]. The general cost threshold $\theta$ is set to 0.2 and the minimum IoU (fallback routine) to 0.3. For all experiments, $N_{init}=1$ and the maximum age parameter is set to $N_{max}=100$ frames (i.e., all tracks that were not successfully matched for more than 100 frames are deleted). With this setting, most occlusions are taken into account while avoiding interference from inter-track occlusions (as shown in Figure 5). As face detector yolov8555https://github.com/ultralytics/ultralytics with a confidence score threshold of 0.4 is used.
@@ 4.2 Evaluated Features
To extract the appearance features, a generic object classification model is used, i.e., a ResNet [45] trained on the ImageNet dataset [52] (resenet18 imported from torchvision models). The appearance features are combined with several different face (biometric) features. The python framework DeepFace [53, 54] provides a wrapper for multiple state-of-the-art face descriptors, i.e., VGG-Face based on the VGG-Very-Deep-16 CNN architecture [39] and evaluated on the Faces in the Wild [55] and the YouTube Faces [56] dataset, Facenet [40] trained on the CASIA-WebFace [57] and the VGGFace2 [44] database, OpenFace [58] trained on the CASIA-WebFace [57] and FaceScrub [59] dataset (based on paper), DeepFace [60] trained on a large collection of Facebook images (i.e., the Social Face Classification (SFC) dataset), DeepID [61] trained on the Faces in the Wild [55] dataset and tested, ArcFace [22, 62], SFace [41] trained on CASIA-WebFace [57] VGGFace2 [44] and MS1MV2 and Dlib [63] trained on the VGG-Face [39] and FaceScrub [59] dataset (plus a large number of images the author scraped from the internet). All these different face descriptors provided by DeepFace are evaluated. DeepFace also provides a 512 dimensional face descriptor for Facenet (i.e., Facenet512) and ArcFace (ArcFaceII) is also evaluated based on a different implementation [^6] https://github.com/mobilesec/arcface-tensorflowlite. In addition, the face descriptor used in [42], SeNet [43], is evaluated.
@@ 4.3 Evaluation Metrics
The Multiple Object Tracking Accuracy (MOTA)[64] is a commonly used metric for evaluating multiple object tracking methods. With MOTA, tracking errors are accumulated on a frame-by-frame basis, i.e., wrongly detected objects (False Positive (FP)), missed objects (False Negative (FN)) and identity switches (IDSWs) are counted for each frame ($t$), summed up and normalized by the total amount of objects in the ground-truth (GT),
@@
An IDSW occurs if the ID assigned in the previous frame ($t-1$) differs from the ID assigned in the current frame ($t$). As MOTA only considers the previous frame, the entire tracking performance of an object during its lifetime is not taken into account. For example, if an object is tracked correctly for 90% of its lifetime but yields the same number of IDSW as an object that is only tracked correctly for 60%, the same MOTA score is achieved. Furthermore, the MOTA score can be dominated by the detector performance (i.e., TP and FP). Since the face detection is fixed for all experiments, only the normalized IDSW is calculated, i.e.,
@@
A metric that focuses on the entire sequence (how long an object is correctly tracked) is the IDF1 [65], i.e.,
@@
is an evaluation metric for assessing MOT trackers, where $DetA_{\alpha}$ and $AssA_{\alpha}$ denote the detection and association accuracy at localization threshold $\alpha$. The localization threshold represents the minimum IoU that a detected BB and a ground-truth BB must reach in order to consider the corresponding ground-truth object as detected (i.e. as TP detection). Thus, HOTA combines all three different aspects for MOT tracker evaluation (i.e., localization, detection and association performance) into a single score. To evaluate the different components individually, HOTA can be decomposed into submetrics. In the considered experimental evaluation, only the association performance is relevant (the detected BBs and locations are fixed for all experiments).
To measure the association performance, the authors in [66] propose the concept of True Positive Associations (TPAs), False Negative Associations (FNAs) and False Positive Associations (FPAs). TPAs are the set of all correctly tracked TPs (correctly detected objects), while FNAs are the set of TPs that are assigned different trIDs for the same ground truth ID (gtID), and FNs (not detected objects). In other words, FNAs represent the amount of intra-person trID switches. On the contrary, FPAs are the set of TPs with the same trID but different gtIDs (i.e., inter-person trID switches), and FPs (wrongly detected objects). Based on the TPAs, FNAs and FPAs the association recall (AssRe), i.e.,
@@
and association precision (AssPr), i.e.,
@@
can be computed. A combination (Jaccard index) of AssRe and AssPr is the association accuracy (AssA), i.e.,
@@
4.4 Experimental Results
In FaceQSORT, the selection of the weighting parameter $\lambda$ (combination of biometric and appearance features) and the association threshold $\theta$ is crucial. Thus, the parameter selection is evaluated in detail along with different face features and a different similarity metric (Euclidean distance instead of cosine similarity).
@@ 4.4.1 Selection of a fixed $\lambda$
Dataset: PLUSFiaQ
Dataset: ChokePoint
Figure 6: The average (across all sequences) AssA and IDF1 scores as well as the normalized amount of IDSW for all evaluated face descriptors and $\lambda$ selections. Recall: for AssA and IDF1, the higher the better, and for IDSW, the lower the better.
@@
ArcFace 0.1 0.6461 0.1 0.7342 0.0 0.0088
DeepFace 0.1 0.6494 0.1 0.7411 0.0 0.0088
DeepID 0.0 0.6303 0.0 0.7277 0.2 0.0083
Dlib 0.0 0.6303 0.0 0.7277 0.4 0.0082
Facenet 0.1 0.6355 0.1 0.7308 0.0 0.0088
Facenet512 0.1 0.6550 0.1 0.7397 0.0 0.0088
OpenFace 0.1 0.6547 0.1 0.7466 0.0 0.0088
SFace 0.1 0.6527 0.1 0.7338 0.0 0.0088
SeNet50 0.2 0.6425 0.2 0.7396 0.2 0.0084
VGG-Face 0.1 0.6515 0.1 0.7405 0.0 0.0088
ArcFaceII 0.1 0.6451 0.1 0.7333 0.0 0.0088
@@
ArcFace 0.2 0.8364 0.1 0.8754 0.3 0.0068
DeepFace 0.6 0.8359 0.3 0.8755 0.0 0.0076
DeepID 0.6 0.8306 0.2 0.8756 0.0 0.0076
Dlib 0.5 0.8353 0.1 0.8752 1.0 0.0052
Facenet 0.3 0.8306 0.0 0.8724 0.0 0.0076
Facenet512 0.1 0.8355 0.1 0.8773 0.1 0.0075
OpenFace 0.2 0.8298 0.2 0.8728 0.0 0.0076
SFace 0.1 0.8404 0.1 0.8783 0.0 0.0076
SeNet50 0.4 0.8325 0.4 0.8740 0.5 0.0069
VGG-Face 0.3 0.8497 0.2 0.8819 0.0 0.0076
ArcFaceII 0.2 0.8396 0.1 0.8777 0.0 0.0076
@@
Figure 7: Average AssPr (across all evaluated sequences).
@@
Figure 6 shows the average (across all sequences) AssA and IDF1 scores as well as the normalized amount of IDSW for all evaluated face descriptors and different $\lambda$ selections. In general, the higher the AssA or IDF1 score or the lower the number of IDSWs, the better the tracking performance. The scores achieved when setting $\lambda$ to 1.0 can be considered as the baseline. In this setting, only face (biometric) features are taken into account (essentially corresponds to StrongSORT with matching cascade). When looking at AssA and IDF1, a general increase in scores can be observed with increasing influence of the appearance feature (i.e. $\lambda\rightarrow 0$). The number of IDSWs, on the other hand, decreases. Thus, the tracking performance increases when $\lambda$ decreases (i.e., with increasing influence of the appearance feature).
@@
The best values achieved (i.e. max. AssA and IDF1 as well as min. IDSW) are shown together with the corresponding $\lambda$ values in tables 2 and 3. In general, significantly higher results are achieved with the ChokePoint dataset (i.e., the PLUSFiaQ dataset is more challenging). With the PLUSFiaQ dataset, the best results (AssA and IDF1) are achieved at $\lambda=0.1$ (except for Dlib, DeepID and SeNet50). In contrast, the best results are achieved with the Chokepoint dataset having $\lambda$ values up to 0.6. One reason for this could be that with the ChokePoint dataset, people look towards the camera when they approach the portal, which facilitates biometric features. The number of IDSWs is usually lowest at $\lambda=0.0$ (i.e., when only appearance features are used). However, it can be stated that the combination of biometric and appearance features improves tracking performance.
@@
In multi-face tracking, inter-person trID switches (i.e., when the trID is switched to an existing trID of a different person) are likely to be more critical than intra-person trID switches (i.e., when a new trID is assigned). As already described, the AssPr score reflects the FPAs (i.e., inter-person trID switches) and is illustrated in Figure 7. If only face (biometric) features are used for association (i.e. $\lambda=1.0$), inter-person identity switches are expected to be limited, as the extracted face (biometric) features should be distinct for different persons. For both analyzed datasets, this can basically be observed for 4/11 evaluated face descriptors (i.e., SFace, ArcFaceII, VGG-Face and Facenet512), with SFace showing the most consistent and best results. However, the AssPr decreases as $\lambda$ decreases. This is reasonable as the association in reducing $\lambda$ is more and more based on generic appearance features and the appearance of two different faces may be similar, while the face (biometric) features are distinct. Considering that the best AssA for these 4 face descriptors is reached at lower $\lambda$ values (i.e., 0.1-0.3 for the ChokePoint and 0.1 for the PLUSFiaQ dataset), the higher AssA is accompanied by a decrease in AssPr (i.e., an increase in inter-person trID switches (FPAs)). In contrast, the AssPr for Dlib and DeepID is much lower for $\lambda=1.0$ than compared to the generic feature baseline (Figure 7 dotted black line). This would indicate weak face descriptors for the considered evaluated sequences.

### v3

@@ 3.1 Computational Complexity
As shown in Algorithm 1, FaceQSORT essentially comprises the matching cascade, the IoU fallback matching, the feature updating and the position prediction. In general, the complexity of finding bipartite matches using the Hungarian algorithm is $O(|\Phi|^{2}\times|\Psi|)$. This is also the worst case with cascade matching. In the considered scenario, it is very likely that the detected face $\Psi$ in frame $t$ were also present in frame $t-1$. Thus, it is reasonable that the complexity can be reduced by the matching cascade (i.e., to $O((\nicefrac{{|\Phi|}}{{x}})^{2}\times\Psi)$). For all unmatched detections after the matching cascade, the complexity of the IoU fallback matching is in the worst case again $O(|\Phi|^{2}\times|\Psi|)$. The complexity for updating the memory for biometric and appearance features ($\Phi_{bio}$ and $\Phi_{app}$) is $O(|\Psi|)$ in each case. Finally, predicting the positions in frame $t+1$ using the Kalman filter has basically a complexity of $O(|\Phi|)$. Thus, the overall complexity of FaceQSORT is,
@@ 4 PLUS Faces in a Queue Dataset
The Paris Lodron University Salzburg Faces in a Queue (PLUSFiaQ) dataset is a new multi-face tracking dataset consisting of a total of seven different sequences. This dataset is made publicly available [^2] https://www.wavelab.at/sources/plusfiaq. An overview of the individual sequences is given in Table 1. In total, the seven sequences comprise 12730 fully annotated frames. This corresponds to about 8 minutes and 30 seconds of video material (25 frames per second). The recorded sequences reflect the assumed scenario, in which several people move towards a gate in a queue (e.g., to enter a sports stadium). Practical use cases for tracking peoples faces as they walk towards a gate are: (i) entrance can be facilitated by registering a face image with the ticket and enabling the person (if correctly recognized) to pass through the gate without additional checks, (ii) potentially unwanted persons (i.e., persons banned from the premises) can be identified on the way to the gate and prevented from entering, (iii) for each ticket presented (which is usually personalized), a set of reference face images can be stored to help identify individuals during the event. Since the time and the area in which a person is present in the visible range of the camera (mounted on the gate) is very limited, the tracking should be highly accurate. In addition, inter-person tracking ID (trID) switches are crucial in the considered scenario. Informed consent was obtained from all participants for all released sequences.
@@ 4.1 Scenario Details
The assumed scenario (i.e., queue in front of a gate) is simulated with 12 different people. These persons move towards a camera equipped gate. On their way to the gate, they form a queue. However, this queue is not very well organized, i.e. people are chatting, eating, moving quickly, walking backwards, pushing or shoving. As soon as a person has entered the gate, they leave the visible area of the camera and thus the scene. To simulate multiple runs, each person goes more than once through the gate. For this reason, the people move in a counterclockwise circle (as depicted in Figure 2). Thus, when a person re-enters the scene, the person first moves away from the gate (camera). This is also illustrated in Figure 3, where the trajectory of the top-left Bounding-Box coordinates from a specific person is shown (i.e., trID 1702, marked by a red circle). At position (a) the person re-enters the scene, then moves away from the gate (b) until he is roughly farthest away at (c) and then moves towards the gate again at (d) and (e). Finally, the person again leaves the scene (goes through the gate) at position (f). This loop is repeated several times.
The only exception is sequence 1, in which only three people are involved, and they move clockwise, i.e. when they re-enter the scene, they move directly towards the gate. This sequence is also the most simple sequence.
@@ 4.2 Annotation Details
The trID is a four-digit number (e.g., xxyy), where the first two digits (xx) represent the personal identifiers and the last two digits (yy) represent a consecutive tracking number (per sequence). A new tracking number is assigned when a person has left the scene (goes through the gate) and re-enters the scene again. For example, the person in figure 3 (a), who is directly in front of the gate, has the personal identifier 18. The current tracking number is 1 (i.e., trID 1801). However, after leaving and re-entering the scene, the tracking number is incremented by one (i.e., trID 1802), as shown in Figure 3 (d). For all non-target class objects (faces) a personal identifier of 99 is assigned. The individual non-target class instances are assigned a dedicated tracking number, independent of how often they have re-entered the scene.
For annotation, the VGG Image Annotator (VIA) [^3] http://www.robots.ox.ac.uk/ vgg/software/via [43] was used. The corresponding exported annotation file (json file) is provided for each sequence. Furthermore, a Ground-Truth (GT) file according to the MOT20 format [44] is provided additionally, i.e. a csv file where each line represents one object instance and contains 9 values:
@@
As a starting point for the annotation, the BBs detected by yolov5444https://github.com/ultralytics/yolov5 and the trIDs predicted by StrongSORT [21] were used. These BBs and trIDs were then refined manually.
@@ 4.3 Tracking Challenges
5 Experimental Setup
FaceQSORT is evaluated based on the proposed PLUSFiaQ (see section 4), the ChokePoint [8] and the MusicVideo [9] dataset. However, the MusicVideo dataset is only utilized when comparing FaceQSORT with state-of-the-art trackers. In the ChokePoint dataset, people moving through a portal (chokepoint) are recorded. The dataset comprises sequences from 18 different runs (i.e., different portal, session or walking direction), with 3 different camera perspectives available for each run (i.e., $18\times 3=54$ available sequences). Of these 54 sequences, only 6 sequences show multiple persons per frame, which is similar to the assumed scenario in this paper. These 6 sequences are therefore used for evaluation. However, no annotations are available for these 6 sequences and are created accordingly. In the ChokePoint dataset, people are reflected in the open glass door as they walk through the portal. These reflections are annotated as non-target classes. The annotations created have the same format as described in section 4 and are made available along with the PLUSFiaQ dataset.
@@ 5.1 Tracker Implementation
FaceQSORT is carefully implemented according to the description in section 3. The parameter $\alpha$ for the EMA feature updating strategy is set to $\alpha=0.9$. The weighting parameter $\beta$ (for weighting the contribution of biometric and appearance similarity compared to spatial distance) is set to 0.98. This gives the biometric and appearance similarity the most weight. $\theta_{pos}$ is set to 9.4877 (the 0.95 quantile of the chi-squared distribution with 4 degrees of freedom). The general cost threshold $\theta$ is set to 0.2 and the minimum IoU (fallback routine) to 0.3. These parameters are set according to [21]. For all experiments, $N_{init}=1$ and the maximum age parameter is set to $N_{max}=100$ frames (i.e., all tracks that were not successfully matched for more than 100 frames are deleted). With this setting, most occlusions are taken into account while avoiding interference from inter-track occlusions (as shown in Figure 5). As face detector yolov8555https://github.com/ultralytics/ultralytics with a confidence score threshold of 0.4 is used.
@@ 5.2 Evaluated Features
To extract the appearance features, a generic object classification model is used, i.e. a ResNet [39] trained on the ImageNet dataset [45] (resenet18 imported from torchvision models). The appearance features are combined with several different face (biometric) features. The python framework DeepFace [46, 47] provides a wrapper for multiple state-of-the-art face recognition models, i.e. VGG-Face (BIO-VG01) based on the VGG-Very-Deep-16 CNN architecture [33] and evaluated on the Faces in the Wild [48] and the YouTube Faces [49] dataset, Facenet [34] (BIO-FN01) trained on the CASIA-WebFace [50] and the VGGFace2 [38] database, OpenFace [51] (BIO-OF01) trained on the CASIA-WebFace [50] and FaceScrub [52] dataset (based on paper), DeepFace [53] (BIO-DF01) trained on a large collection of Facebook images (i.e., the Social Face Classification (SFC) dataset), DeepID [54] (BIO-DI01) trained on the Faces in the Wild [48] dataset and tested, ArcFace [26, 55] (BIO-AF01), SFace [35] (BIO-SF01) trained on CASIA-WebFace [50] VGGFace2 [38] and MS1MV2 and Dlib [56] (BIO-DL01) trained on the VGG-Face [33] and FaceScrub [52] dataset (plus a large number of images the author scraped from the internet). All these different face recognition models provided by DeepFace are evaluated. DeepFace also provides a 512 dimensional face descriptor for Facenet (BIO-FN02) and ArcFace (BIO-AF02) is also evaluated using different weights [^6] https://www.digidow.eu/f/datasets/arcface-tensorflowlite/model.tflite. In addition, the face recognition model used in [36], SeNet [37] (BIO-SN01), is evaluated.
@@ 5.3 Evaluation Metrics
The Multiple Object Tracking Accuracy (MOTA)[57] is a commonly used metric for evaluating multiple object tracking methods. With MOTA, tracking errors are accumulated on a frame-by-frame basis, i.e. wrongly detected objects (False Positive (FP)), missed objects (False Negative (FN)) and identity switches (IDSWs) are counted for each frame ($t$), summed up and normalized by the total amount of objects in the ground-truth (GT),
@@
An IDSW occurs if the ID assigned in the previous frame ($t-1$) differs from the ID assigned in the current frame ($t$). As MOTA only considers the previous frame, the entire tracking performance of an object during its lifetime is not taken into account. For example, if an object is tracked correctly for 90% of its lifetime but yields the same number of IDSW as an object that is only tracked correctly for 60%, the same MOTA score is achieved. Furthermore, the MOTA score can be dominated by the detector performance (i.e., TP and FP). Since the face detection is fixed for all experiments, only the normalized IDSW is calculated, i.e.
@@
A metric that focuses on the entire sequence (how long an object is correctly tracked) is the IDF1 [58], i.e.
@@
is an evaluation metric for assessing MOT trackers, where $DetA_{\alpha}$ and $AssA_{\alpha}$ denote the detection and association accuracy at localization threshold $\alpha$. The localization threshold represents the minimum IoU that a detected BB and a ground-truth BB must reach in order to consider the corresponding ground-truth object as detected (i.e., as TP detection). Thus, HOTA combines all three different aspects for MOT tracker evaluation (i.e., localization, detection and association performance) into a single score. To evaluate the different components individually, HOTA can be decomposed into submetrics. In the considered experimental evaluation, only the association performance is relevant (the detected BBs and locations are fixed for all experiments).
To measure the association performance, the authors in [59] propose the concept of True Positive Associations (TPAs), False Negative Associations (FNAs) and False Positive Associations (FPAs). TPAs are the set of all correctly tracked TPs (correctly detected objects), while FNAs are the set of TPs that are assigned different trIDs for the same ground truth ID (gtID), and FNs (not detected objects). In other words, FNAs represent the amount of intra-person trID switches. On the contrary, FPAs are the set of TPs with the same trID but different gtIDs (i.e., inter-person trID switches), and FPs (wrongly detected objects). Based on the TPAs, FNAs and FPAs the association recall (AssRe), i.e.
@@
and association precision (AssPr), i.e.
@@
can be computed. A combination (Jaccard index) of AssRe and AssPr is the association accuracy (AssA), i.e.
@@
The code provided in [60] is used to calculate all evaluation metrics, with the average value (across all sequences) usually being reported.
6 Experimental Evaluation
A comprehensive experimental evaluation is conducted, including a parameter evaluation, a comparison with the state-of-the-art and an ablation study. Using Euclidean distance as similarity metric for association instead of cosine similarity lead to a different scaling of the matching scores. However, the tracking performance does not change significantly. Thus (and due to space constraints), these results are not reported in this work. Interested readers will find these results in the supplementary material together with the complete results, from which all figures and tables are generated. In all tables, the best results are marked in bold.
6.1 Parameter Evaluation
In FaceQSORT, the selection of the weighting parameter $\lambda$ (combination of biometric and appearance features) and the association threshold $\theta$ is crucial. Thus, the parameter selection is evaluated in detail along with biometric features extracted by different face recognition models.
@@ 6.1.1 Selection of a fixed $\lambda$
Figure 6: The average (across all sequences) AssA and IDF1 scores as well as the normalized amount of IDSW for all evaluated face recognition models and $\lambda$ selections. Recall: for AssA and IDF1, the higher the better, and for IDSW, the lower the better.
@@
BIO-AF01 0.1 0.6461 0.1 0.7342 0.0 0.0088
BIO-AF02 0.1 0.6451 0.1 0.7333 0.0 0.0088
BIO-DF01 0.1 0.6494 0.1 0.7411 0.0 0.0088
BIO-DI01 0.0 0.6303 0.0 0.7277 0.2 0.0083
BIO-DL01 0.0 0.6303 0.0 0.7277 0.4 0.0082
BIO-FN01 0.1 0.6355 0.1 0.7308 0.0 0.0088
BIO-FN02 0.1 0.6550 0.1 0.7397 0.0 0.0088
BIO-OF01 0.1 0.6547 0.1 0.7466 0.0 0.0088
BIO-SF01 0.1 0.6527 0.1 0.7338 0.0 0.0088
BIO-SN01 0.2 0.6425 0.2 0.7396 0.2 0.0084
BIO-VG01 0.1 0.6515 0.1 0.7405 0.0 0.0088
@@
BIO-AF01 0.2 0.8364 0.1 0.8754 0.3 0.0068
BIO-AF02 0.2 0.8396 0.1 0.8777 0.0 0.0076
BIO-DF01 0.6 0.8359 0.3 0.8755 0.0 0.0076
BIO-DI01 0.6 0.8306 0.2 0.8756 0.0 0.0076
BIO-DL01 0.5 0.8353 0.1 0.8752 1.0 0.0052
BIO-FN01 0.3 0.8306 0.0 0.8724 0.0 0.0076
BIO-FN02 0.1 0.8355 0.1 0.8773 0.1 0.0075
BIO-OF01 0.2 0.8298 0.2 0.8728 0.0 0.0076
BIO-SF01 0.1 0.8404 0.1 0.8783 0.0 0.0076
BIO-SN01 0.4 0.8325 0.4 0.8740 0.5 0.0069
BIO-VG01 0.3 0.8497 0.2 0.8819 0.0 0.0076
@@
Figure 7: Average AssPr (across all evaluated sequences) for different $\lambda$ selections.
@@
Figure 6 shows the average (across all sequences) AssA and IDF1 scores as well as the normalized amount of IDSW for all evaluated face recognition models and different $\lambda$ selections. In general, the higher the AssA or IDF1 score or the lower the number of IDSWs, the better the tracking performance. The scores achieved when setting $\lambda$ to 1.0 can be considered as the baseline. In this setting, only biometric features are taken into account. When looking at AssA and IDF1, a general increase in scores can be observed with increasing influence of the appearance feature (i.e., $\lambda\rightarrow 0$). The number of IDSWs, on the other hand, decreases. Thus, the tracking performance increases when $\lambda$ decreases (i.e., with increasing influence of the appearance feature).
@@
The best values achieved (i.e., max. AssA and IDF1 as well as min. IDSW) are shown together with the corresponding $\lambda$ values in tables 2 and 3. In general, significantly higher results are achieved with the ChokePoint dataset (i.e., the PLUSFiaQ dataset is more challenging). With the PLUSFiaQ dataset, the best results (AssA and IDF1) are achieved at $\lambda=0.1$ (except for BIO-DI01, BIO-DL01 and BIO-SN01). In contrast, the best results are achieved with the Chokepoint dataset having $\lambda$ values up to 0.6. One reason for this could be that with the ChokePoint dataset, people look towards the camera when they approach the portal, which facilitates the usage of biometric features. The number of IDSWs is usually lowest at $\lambda=0.0$ (i.e., when only appearance features are used). However, it can be stated that the combination of biometric and appearance features improves tracking performance for both datasets.
@@
In the considered scenario, inter-person trID switches (i.e., when the trID is switched to an existing trID of a different person) are more critical than intra-person trID switches (i.e., when a new trID is assigned). As already described, the AssPr score reflects the FPAs (i.e., inter-person trID switches) and is illustrated in Figure 7. If only biometric features are used for association (i.e., $\lambda=1.0$), inter-person trID switches are expected to be limited, as the extracted biometric features should be distinct for different identities. For both analyzed datasets, this can basically be observed for 4/11 evaluated face recognition models (i.e., BIO-SF01, BIO-AF02, BIO-FN02 and BIO-VG01), with BIO-SF01 features showing the most consistent and best results. However, the AssPr decreases as $\lambda$ decreases. This is reasonable as the association in reducing $\lambda$ is more and more based on generic appearance features and the appearance of two different identities may be similar, while the biometric features are distinct. Considering that the best AssA for these 4 face recognition models is reached at lower $\lambda$ values (i.e., 0.1 to 0.3 for the ChokePoint and 0.1 for the PLUSFiaQ dataset), the higher AssA is accompanied by a decrease in AssPr (i.e., an increase in inter-person trID switches (FPAs)). In contrast, the AssPr for BIO-DL01 and BIO-DI01 is much lower for $\lambda=1.0$ than compared to the appearance features baseline (Figure 7 dotted black line). This would indicate a weak distinctiveness of these biometric features for the considered scenario.